The pressure exerted by 12g of an ideal gas at temperature is increased by 10 degree Celsius at the same volume ,the pressurest increases by 10% calculate the temperature t and and volume 
Van.molecular weight of the gas =120

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear stdent 
Given P = 1 atm, w = 12 g; T = (t + 273)K; V = V litre
 If T = t + 10 + 273 = t + 283 K; V = V litre,
P = 1 + 10/100 = 1.1 atm
Using gas equation, PV = w/m RT
Case I. 1 * V = 12/m R (t + 273) ….(1)
Case II. 1.1 * V = 12/m R (t + 283) ….(2)
From (1) and (2), t = - 137oC or t = 100 K
Also from (1), on substituting t and m (120), V = 0.82 litre
Hope this helps